OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 29

OpenStudy (amistre64):

lets find the common difference an = a1 + d(n-1) a6 = 4 + d(5) = 29; solve for d

OpenStudy (amistre64):

once we know d, its just using the general formula again for n=32 a32 = a1 + d(32-1)

OpenStudy (anonymous):

why are there 2 equal signs in the first equation ?

OpenStudy (anonymous):

it means a=b=c which means a=b b=c a=c

OpenStudy (anonymous):

same as 1+3=4=5-1

OpenStudy (anonymous):

1+3=4 4=5-1 5-1=1+3